OpenStudy (anonymous):

TRUE OR FALSE...The converse of the statement “If x – 4 = 17, then x = 21”

OpenStudy (anonymous):

Do you think x can be any other value?

OpenStudy (anonymous):

lol no it should be a "Y" right

OpenStudy (anonymous):

"Y" as in true?

OpenStudy (anonymous):

I was answering your question!....but the answer to my question above is FALSE?

OpenStudy (anonymous):

It doesnt matter what letter x is? Anyways if you take x- 4 = 17, and add 4 to each side, x- 4 + 4 = 17 + 4 x = 17 + 4 = ?

OpenStudy (anonymous):

oooooooh...so it is TRUE,, thanks

OpenStudy (anonymous):

No probs